Campbell Biology, 10edition

Answer all of the following questions

Chapter 14   Mendel and the Gene Idea

a) What is a monohybrid cross and dihybrid cross? How do the number of characters being studied and whether the trait is homozygous or heterozygous play a role?

b) Calculate the different gametes could be produced through independent assortment by an individual with the genotype KkLlMmDdEE?

c) What are the gametes that could be produced from a parent that is AaBb.

d) How many different gametes can a parent produce if it has 4 different alleles for a single gene? What is the maximum number of possible different gametes that this parent could produce?

e) Describe the progeny of one parent that is dominate and one parent is recessive. Give an examples of what offspring would be expected over three generations.

f) Calculate the ratio and percentage of a progeny produced by the cross: KkLl× KKll.

g) What is the expect ratio of all phenotypes from a cross between Ggss x GgSS?

h) Calculate the F2 generation rations from a monohybrid cross; --- a dihybrid cross.

i)       What is phenylketonuria and how is it related to genetics?   What are the expected offspring between two carriers? What is likelihood of having a normal son?

j) What is incomplete dominance and how is it physically displayed in organisms?

Give an example of incomplete dominance and calculate the ratio of expected offspring from various parent combinations.

k) What is pleiotropy and how is it physically displayed in organisms?

Give an example of pleiotropy and calculate the ratio of expected offspring from various parent combinations.

l) What is incomplete dominance and how is it physically displayed in organisms?

Give an example of incomplete dominance when two incomplete dominant traits are involved and calculate the ratio of expected offspring from various parent combinations.

m) Discuss the concept of ABO blood types in humans.   Show your understanding by being able to calculate the offspring possibilities from any combination of two given parents (the blood types of the parents is given).
Answer #1

a) Monohybrid cross is a cross between two diploid individuals, considering only one characteristic for e.g. "AA" crossed to "aa".

Dihybrid cross is a cross when two characteristics/factors/genes are considered for e.g. "AABB" crossed to "aabb".

Whatever the number of traits is there, all alleles of a gene assort independently; and all genes are segregated during gamete formation.

So, for a monohybrid cross "Aa" genotype will be segregated into two gametes, one gamete will have "A" and the other will have "a" allele. A homozygous individual will produce gametes having same alleles will a heterozygous individuals will have different alleles in different gametes. If more than one traits are there, then each gamete will have one allele of each trait.

b) The individual is "KkLlMmDdEE".

Total number of gametes = 2 (allele number -gene number) = 2 (9-5) = 2(4) = 16 gametes.

c) A parent that is AaBb will produce four gametes: AB, Ab, aB, and ab.

d) There are 4 alleles for one gene. So, no. of gametes = 2 (4-1) = 2(3) = 8 = This is the maximum number of different alleles this parent can produce.